# Improving Bilinear RNNs with Closed-loop Control

### v4

@@ 2.1 Linear RNNs
Unlike autoregressive Transformers that store all contextual information in KV cache, linear RNNs compress highly abstract knowledge into a fixed-size state for generalization, structurally resembling energy-based models [56], i.e., Hopfield networks [65, 31] and neural Hebbian learning systems [18]. Early models like Linformer [98], S4 [41] and RetNet [91] lack effective, data-dependent memory control, resulting in inferior performance to softmax attention. Later models like Mamba [38] and GLA [107] address this by introducing a dynamic, projection-based gating, yielding substantial improvements. Formally, these models are linear register systems with key-value associative memory, where the memory is written by the forgetting/input gates ($\bm{\alpha}$, $\bm{\beta}$) and retrieved via query-based read:
@@
Due to differences in theoretical foundations and development trajectories, linear RNNs have evolved into two main implementation paradigms: (i) linear attentions (LAs) [112, 79, 22, 107, 105, 5] and (ii) state space models (SSMs) [72, 38, 39], we summarize their key distinctions in Table 2:
@@
i) State size: SSMs like Mamba2 [26] adopt a fixed state expand dimension of 128, resulting in a state size of 256D. Whereas in LAs, the state size is determined by the dimension of key/value heads (typically 64). In practice, Mamba2 offers a higher capacity [20], which underpins its advantages in retrieval tasks and hybrid architectures [88, 89].
ii) Parameter composition: SSMs like Mamba2 resembles a multi-value attention mechanism [26], where the input projection $\bar{\bm{B}}$ ($\bm{k}$) and output projection $\bar{\bm{C}}$ ($\bm{q}$) are shared across all value heads $\bm{u}$} ($\bm{v}$). Additionally, state space models omit the feedforward network and instead double both the input dimension and the number of layers to increase model capacity. In this paper, we empirically follow the linear attention design but keep the head dimension at 256 to match the state size of Mamba2. A detailed architectural ablation is in §4.2.
Chunk-wise Parallel  Although Linear RNNs achieve a favorable pretraining time complexity of $\mathcal{O}(LD^{2})$, they often slower than softmax attention with $\mathcal{O}(L^{2}D)$ complexity on shorter sequences. This is mainly due to the fact that current hardware is highly optimized for matmul operations, which limits the efficiency of linear recurrence, necessitating additional training optimizations. S4 [41] introduces a complex diagonal-plus-low-rank design using the Cauchy kernel, which is later simplified in DSS [43] and S4D [40] through complex and real diagonal approximations to employ Vandermonde-based convolutions. Other models like S5 [84] and Mamba [38] leverage the Blelloch scan algorithm [15] to cache intermediate results and speed up recurrent computation. Recent methods inspired by FlashAttention [25], including Lightning-Attns [77, 78, 58], GLA [107], and Mamba2 [26], introduce inter-chunk recurrence combined with intra-chunk parallelism to fully utilize matrix compute throughput. A basic formulation using chunk size $C$ can be expressed as:
@@ 2.2 Bilinear RNNs and Beyond
In a neural memory perspective [34], effective memory management remains a central challenge. Unlike Hebbian learning rule [18], which relies on reinforcement-based memory updates, Delta learning rule [75, 101] focuses on supervised memory control and has been extensively explored in various works, such as fast weight programs [50, 81] and Meta-learning [68, 67]. (Gated-)DeltaNet [108, 105] employs it as a memory correction $\bm{v}_{t}^{\text{new}}=\bm{v}_{t}-\bm{S}_{t-1}\bm{k}_{t}$, and based on Eq. 2, introduces an efficient chunk-wise parallel algorithm for hardware-efficient training in sequence modeling:
@@
These models are no longer a linear key-value memory register as in Eq. 1; instead, the interaction between the state $\bm{S}$ and the input $\bm{k}$ introduces a bilinear term $\bm{Sk}$, this resembles a affine bilinear system [^1] Bilinear systems are linear with respect to state and input individually, but nonlinear overall due to the product term (e.g., $\bm{Sk}$). They are regarded as a special class of nonlinear systems that preserve controllability. [17, 113, 99, 70]. Accordingly, models that are similar to the updating rule in Eq. 3 can be classified as Bilinear RNNs, and their key strengths are summarized as follows:
@@
As shown in Fig. 1, the Householder transform $(\bm{I}-\beta_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal})$ generated by the Delta rule defines a mirror transform, effectively reflecting stored memories across a hyper-plane orthogonal to $\bm{k}_{t}$. Given a memory $\bm{S}\in\mathbb{R}^{D\times D}$ with at most $D$ orthogonal memory patterns, when the sequence length $t>D$, to avoid memory conflicts, this reflection attenuates components of stored memories $\{\bm{S}_{t-1}^{i}\}_{i=1}^{D}$ in non-orthogonal directions based on factor $\beta_{t}$. This mechanism implicitly enforces orthogonal memory management, enabling the model to preserve $D$ distinguishable historical memories over time. In this paper, Comba modifies the state transition to $(\alpha_{t}-\beta_{t}\bm{k}\bm{k}^{\intercal})$, which is a scalar-plus-low-rank form (SPLR), enabling more flexible supervision. To some extent, this process can be seen as a Schmidt orthogonalization [57] or a rotation operation [86] on the KV cache.
@@
Linear RNNs generally approximate the dense state transition matrix [66, 46] with a diagonal matrix or scalar [40, 26], significantly reducing computational overhead but at the cost of expressiveness [1]. While the additional low-rank terms of the state transition in Eq. 3 improve the model’s expressiveness while preserving tractability for efficient parallelization.
Building on this, Gated-DeltaNet [105] introduces a global scalar forgetting gate on the state, and Delta-Product [83] explore a multi-step Householder transform $\prod_{j=1}^{n}(\bm{I}-\beta_{t,j}\bm{k}_{t,j}\bm{k}_{t,j}^{\top})$, enabling a smooth interpolation between purely diagonal and fully dense transitions. RWKV7 [73] improves the IPLR form to diagonal-plus-low-rank (DPLR) form, which aligns with insights from HiPPO theory [39, 42], which shows that all orthogonal polynomial projection matrices can be decomposed into DPLR components. Our proposed Comba adopts a scalar-plus-low-rank (SPLR) form. Empirically, we find that the scalar is sufficiently expressive (similar to the empirical simplification from Mamba1 to Mamba2) and offers significant pretraining acceleration over RWKV-7.
From another perspective, models such as TTT [90] separate the model weights into inner and outer components, updating the inner memory directly via stochastic gradient descent (SGD) [16].
@@
where the state $\bm{S}$ can be parameterized by a matrix or a two-layer MLP to increase memory capacity. When mini-batch $B=1$ and ignoring normalization operation, this reduces to the update rule in Eq. 3, and when adopting MLP-based deep memory, the model transitions into a modern form of Nonlinear RNN. Subsequently, Titans [7] introduces a data-dependent state decay and momentum, while models like MIRAS [6] and Lattice [52] enhance $\bm{S}$ using nonlinear higher-order optimization, e.g., $\operatorname{sign}$ function or high-order derivatives. However, there is no free lunch: these models rely on mini-batch gradient descent approach to compensate for the sequence-level parallelism barrier introduced by their structural complexity, but empirical results show that a mini-batch of 1 (similar to Eq. 3, which is a first-order approach) remains optimal [90], especially in language modeling. Comba retains the bilinear form and offers an efficient chunk-wise parallel optimization.
@@ 3 Bilinear RNNs in Closed-loop Control
Unlike perspectives from neural memory or optimization, this work revisits Bilinear RNNs through the lens of control theory [23, 24, 87]. As shown in Table 3, linear RNNs such as Mamba2 and GLA are generally viewed, to some extent, as open-loop control systems, where the output does not provide feedback to influence the control behavior. In contrast, another class of systems, known as closed-loop control systems [45], incorporates negative feedback to enhance the adaptability of the system and allows it to handle more complex dynamic tasks. According to Wikipedia, a closed-loop controller should use feedback to control states or outputs of a dynamical system. So in this paper, Comba adopts a two-stage feedback strategy: the input information $\bm{v}_{t}$ is first corrected via state-based feedback $\bm{P}_{t}(\cdot)$, and the output is similarly refined using the same feedback mechanism. Compared to directly output correction methods, e.g., use $\bm{q}_{t}$ to compute $\bm{v}_{t}^{\text{new}}$, this approach is generally considered more robust and better suited for parallel computation (as it only modifies $\bm{q}_{t}$ at the output stage without involving it in state updates). From this perspective, models such as TTT, DeltaNet, and RWKV-7 incorporate only the first-stage state feedback correction.
@@
Similar to the optimizer perspective, the feedback $\bm{P}_{t}(\cdot)$ in closed-loop control can be either linear and first-order, or nonlinear and higher-order. However, two major challenges arise: (i) adopting nonlinear or high-order optimization techniques, as in TTT, Lattice, or MIRAS, will hinder chunk-wise parallelism; and (ii) recurrent models suffer from the well-known issue of exponential gradient explosion [74] during training. If we follow DeltaNet using first-order feedback but initialize a new vector [106] to interact with the state, it becomes difficult to ensure that the spectral radius of the state transition matrix remains below one. Therefore, considering these factors, Comba follows previous models using the $\bm{k}$ vector to interact with the state $\bm{S}$, while introducing a special treatment of feedback strength detailed in the following sections. For a standard Comba with forget gate $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}$}, state feedback factor $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\beta}_{t}$}, input gate $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}$}, and output feedback factor $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}d$}:
@@
In practice, the scalar gating will be limited to the interval $(0,1)$, and $\|\bm{k}_{t}\|=1$ (L2 Norm). DeltaNet formulates the state transition in an IPLR structure, while RWKV-7 improves expressiveness via a DPLR extension with LoRA-style [47] diagonal initialization, i.e., $\operatorname{diag}(\bm{w}_{t})-\beta_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal}$. However, our results indicate that the low-rank form will impair model capacity. Moreover, recent efforts [12, 55, 11, 69] aim to distill Transformers into recurrent structures to leverage prior knowledge of large-scale pretrained Transformers, where a key design principle is to minimize additional parameters. To this end, Comba adopts an SPLR structure that only introduces a data-dependent scalar, achieving superior empirical performance compared to RWKV-7 (as implemented in FLA [109]), along with a 2 $\times$ pretraining acceleration (the memory usage is only half). Recent works [83, 36] have also improved IPLR structures by extending their eigenvalues into the negative domain, e.g., $\bm{I}-2\beta_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal}$, to enhance the model’s state-tracking capability. Notably, SPLR naturally admits negative eigenvalues. For fair comparison, in Table 4, we introduce a variant named Comba-$\operatorname{iplr}$. However, such models with global state decay tend to overfit, and the SPLR structure remains optimal and is more structurally aligned with control theory.
@@
Comba introduces additional output feedback, from an optimization perspective, this is equivalent to incorporating a similarity optimization objective $\left\langle\bm{q}_{t},d\bm{k}_{t}\right\rangle$ with factor $d$. In a neural memory perspective, $\bm{k}$ ensures that memory $\bm{v}$ is stored as clearly as possible, enabling precise querying by $\bm{q}$. This optimization objective directly facilitates this process, and significantly reduces the model’s perplexity, thereby enhancing performance (§4.2). Empirically, initializing $d$ to 0.02 improves performance for some smaller models (e.g., 340M), enabling gradual learning of the similarity between $\bm{q}$ and $\bm{k}$. For larger models (e.g., 1.3B, 2.7B), initializing $d$ to 1 leads to the greatest performance [^2] We find that $d=1$ yields better performance in most cases, which is the default configuration for Comba.. While prior research has largely focused on improving gating mechanisms or optimizing state updates via key-value operations, few models have explicitly modified the output pathway (i.e., the query @ State). We think there appears to be a potential connection between Comba and MesaNet [96], as MesaNet also performs query correction in the output stage by solving a closed-form recursive least squares problem [10], resulting in $\bm{q}_{t}=\left(\bm{H}_{t}+\bm{\Lambda}\right)^{-1}\bm{q}_{t}$. Notably, the correction in Comba can be implemented with a single line of code $\bm{q}_{t}-d\bm{k}_{t}$ and is applicable to nearly all RNN variants. In App. A.2, we also provide an alternative explanation, namely that $d$ is equivalent to the residual matrix $\bm{D}$ in Mamba [38].
@@
Model Forget Gate $\bm{\alpha}_{t}$ Range Input Gate $\bm{\beta}_{t}$ Range GLA [107] $\operatorname{sigmoid}(\bm{W}_{1}\bm{W}_{2}\bm{x}_{t})^{\frac{1}{\tau}}\bm{1}^{\intercal}$ (0, 1) N/A 1 Mamba2 [26] $\exp\left(-a\operatorname{softplus}\left(\bm{W}_{\alpha}\bm{x}_{t}+c\right)\right)$ $\sim$ 1 $\operatorname{softplus}\left(\bm{W}_{\alpha}\bm{x}_{t}+c\right)$ $\sim 0$ MetaLA [22] $\operatorname{sigmoid}(\bm{W}_{\alpha}\bm{x}_{t})\bm{1}^{\intercal}$ (0, 1) $\bm{1}-\operatorname{sigmoid}(\bm{W}_{\alpha}\bm{x}_{t})$ (0, 1) Comba (ours) $\exp\left(-a\operatorname{softplus}\left(\bm{W}_{\alpha}\bm{x}_{t}+c\right)\right)$ $\sim$ 1 $\operatorname{sigmoid}(\bm{W}_{\beta}\bm{x}_{t})$ $\tilde{\beta}_{t}$ << $\beta_{t}$ $\in$ (0, 1)

### v5

@@ 2.1 Linear RNNs
Unlike autoregressive Transformers that store all contextual information in KV cache, linear RNNs compress highly abstract knowledge into a fixed-size state for generalization, structurally resembling energy-based models [lecun2006tutorial], i.e., Hopfield networks [mceliece1987capacity, farhat1985optical] and neural Hebbian learning systems [chakraverty2019hebbian]. Early models like Linformer [wang2020linformer], S4 [gu2021efficiently] and RetNet [sun2023retentive] lack effective, data-dependent memory control, resulting in inferior performance to softmax attention. Later models like Mamba [gu2023mamba] and GLA [yang2023gated] address this by introducing a dynamic, projection-based gating, yielding substantial improvements. Formally, these models are linear register systems with key-value associative memory, where the memory is written by the forgetting/input gates ($\bm{\alpha}$, $\bm{\beta}$) and retrieved via query-based read:
@@
Due to differences in theoretical foundations and development trajectories, linear RNNs have evolved into two main implementation paradigms: (i) linear attentions (LAs) [zhang2024gated, qin2024hgrn2, chou2024metala, yang2023gated, yang2024gated, beck2024xlstm] and (ii) state space models (SSMs) [patro2024mamba, gu2023mamba, gu2020hippo], we summarize their key distinctions in Table 2:
@@
i) State size: SSMs like Mamba2 [dao2024transformers] adopt a fixed state expand dimension of 128, resulting in a state size of 256D. Whereas in LAs, the state size is determined by the dimension of key/value heads (typically 64). In practice, Mamba2 offers a higher capacity [chen2024stuffed], which underpins its advantages in retrieval tasks and hybrid architectures [sun2025linear, sun2024hunyuan].
ii) Parameter composition: SSMs like Mamba2 resembles a multi-value attention mechanism [dao2024transformers], where the input projection $\bar{\bm{B}}$ ($\bm{k}$) and output projection $\bar{\bm{C}}$ ($\bm{q}$) are shared across all value heads $\bm{u}$} ($\bm{v}$). Additionally, state space models omit the feedforward network and instead double both the input dimension and the number of layers to increase model capacity. In this paper, we empirically follow the linear attention design but keep the head dimension at 256 to match the state size of Mamba2. A detailed architectural ablation is in §4.2.
Chunk-wise Parallel  Although Linear RNNs achieve a favorable pretraining time complexity of $\mathcal{O}(LD^{2})$, they often slower than softmax attention with $\mathcal{O}(L^{2}D)$ complexity on shorter sequences. This is mainly due to the fact that current hardware is highly optimized for matmul operations, which limits the efficiency of linear recurrence, necessitating additional training optimizations. S4 [gu2021efficiently] introduces a complex diagonal-plus-low-rank design using the Cauchy kernel, which is later simplified in DSS [gupta2022diagonal] and S4D [gu2022parameterization] through complex and real diagonal approximations to employ Vandermonde-based convolutions. Other models like S5 [smith2022simplified] and Mamba [gu2023mamba] leverage the Blelloch scan algorithm [blelloch1990prefix] to cache intermediate results and speed up recurrent computation. Recent methods inspired by FlashAttention [dao2022flashattention], including Lightning-Attns [qin2024lightning, qin2024various, li2025minimax], GLA [yang2023gated], and Mamba2 [dao2024transformers], introduce inter-chunk recurrence combined with intra-chunk parallelism to fully utilize matrix compute throughput. A basic formulation using chunk size $C$ can be expressed as:
@@ 2.2 Bilinear RNNs and Beyond
In a neural memory perspective [gershman2025key], effective memory management remains a central challenge. Unlike Hebbian learning rule [chakraverty2019hebbian], which relies on reinforcement-based memory updates, Delta learning rule [prados1989neural, widrow1988adaptive] focuses on supervised memory control and has been extensively explored in various works, such as fast weight programs [irie2021going, schlag2021linear] and Meta-learning [munkhdalai2017neural, munkhdalai2019metalearned]. (Gated-)DeltaNet [yang2024parallelizing, yang2024gated] employs it as a memory correction $\bm{v}_{t}^{\text{new}}=\bm{v}_{t}-\bm{S}_{t-1}\bm{k}_{t}$, and based on Eq. 2, introduces an efficient chunk-wise parallel algorithm for hardware-efficient training in sequence modeling:
@@
These models are no longer a linear key-value memory register as in Eq. 1; instead, the interaction between the state $\bm{S}$ and the input $\bm{k}$ introduces a bilinear term $\bm{Sk}$, this resembles a affine bilinear system [^1] Bilinear systems are linear with respect to state and input individually, but nonlinear overall due to the product term (e.g., $\bm{Sk}$). They are regarded as a special class of nonlinear systems that preserve controllability. [bruni1974bilinear, zhao2016gramian, wang2023expectation, pardalos2010optimization]. Accordingly, models that are similar to the updating rule in Eq. 3 can be classified as Bilinear RNNs, and their key strengths are summarized as follows:
@@
As shown in Fig. 1, the Householder transform $(\bm{I}-\beta_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal})$ generated by the Delta rule defines a mirror transform, effectively reflecting stored memories across a hyper-plane orthogonal to $\bm{k}_{t}$. Given a memory $\bm{S}\in\mathbb{R}^{D\times D}$ with at most $D$ orthogonal memory patterns, when the sequence length $t>D$, to avoid memory conflicts, this reflection attenuates components of stored memories $\{\bm{S}_{t-1}^{i}\}_{i=1}^{D}$ in non-orthogonal directions based on factor $\beta_{t}$. This mechanism implicitly enforces orthogonal memory management, enabling the model to preserve $D$ distinguishable historical memories over time. In this paper, Comba modifies the state transition to $(\alpha_{t}-\beta_{t}\bm{k}\bm{k}^{\intercal})$, which is a scalar-plus-low-rank form (SPLR), enabling more flexible supervision. To some extent, this process can be seen as a Schmidt orthogonalization [leon2013gram] or a rotation operation [su2024roformer] on the KV cache.
@@
Linear RNNs generally approximate the dense state transition matrix [medsker2001recurrent, hochreiter1997long] with a diagonal matrix or scalar [gu2022parameterization, dao2024transformers], significantly reducing computational overhead but at the cost of expressiveness [amos2023never]. While the additional low-rank terms of the state transition in Eq. 3 improve the model’s expressiveness while preserving tractability for efficient parallelization.
Building on this, Gated-DeltaNet [yang2024gated] introduces a global scalar forgetting gate on the state, and Delta-Product [siems2025deltaproduct] explore a multi-step Householder transform $\prod_{j=1}^{n}(\bm{I}-\beta_{t,j}\bm{k}_{t,j}\bm{k}_{t,j}^{\top})$, enabling a smooth interpolation between purely diagonal and fully dense transitions. RWKV7 [peng2025rwkv] improves the IPLR form to diagonal-plus-low-rank (DPLR) form, which aligns with insights from HiPPO theory [gu2020hippo, gu2022train], which shows that all orthogonal polynomial projection matrices can be decomposed into DPLR components. Our proposed Comba adopts a scalar-plus-low-rank (SPLR) form. Empirically, we find that the scalar is sufficiently expressive (similar to the empirical simplification from Mamba1 to Mamba2) and offers significant pretraining acceleration over RWKV-7.
From another perspective, models such as TTT [sun2024learning] separate the model weights into inner and outer components, updating the inner memory directly via stochastic gradient descent (SGD) [bottou2010large].
@@
where the state $\bm{S}$ can be parameterized by a matrix or a two-layer MLP to increase memory capacity. When mini-batch $B=1$ and ignoring normalization operation, this reduces to the update rule in Eq. 3, and when adopting MLP-based deep memory, the model transitions into a modern form of Nonlinear RNN. Subsequently, Titans [behrouz2024titans] introduces a data-dependent state decay and momentum, while models like MIRAS [behrouz2025s] and Lattice [karami2025lattice] enhance $\bm{S}$ using nonlinear higher-order optimization, e.g., $\operatorname{sign}$ function or high-order derivatives. However, there is no free lunch: these models rely on mini-batch gradient descent approach to compensate for the sequence-level parallelism barrier introduced by their structural complexity, but empirical results show that a mini-batch of 1 (similar to Eq. 3, which is a first-order approach) remains optimal [sun2024learning], especially in language modeling. Comba retains the bilinear form and offers an efficient chunk-wise parallel optimization.
@@ 3 Bilinear RNNs in Closed-loop Control
Unlike perspectives from neural memory or optimization, this work revisits Bilinear RNNs through the lens of control theory [collins1993open, collins1995age, sun2016open]. As shown in Table 3, linear RNNs such as Mamba2 and GLA are generally viewed, to some extent, as open-loop control systems, where the output does not provide feedback to influence the control behavior. In contrast, another class of systems, known as closed-loop control systems [hjalmarsson2005experiment], incorporates negative feedback to enhance the adaptability of the system and allows it to handle more complex dynamic tasks. According to Wikipedia, a closed-loop controller should use feedback to control states or outputs of a dynamical system. So in this paper, Comba adopts a two-stage feedback strategy: the input information $\bm{v}_{t}$ is first corrected via state-based feedback $\bm{P}_{t}(\cdot)$, and the output is similarly refined using the same feedback mechanism. Compared to directly output correction methods, e.g., use $\bm{q}_{t}$ to compute $\bm{v}_{t}^{\text{new}}$, this approach is generally considered more robust and better suited for parallel computation (as it only modifies $\bm{q}_{t}$ at the output stage without involving it in state updates). From this perspective, models such as TTT, DeltaNet, and RWKV-7 incorporate only the first-stage state feedback correction.
@@
Similar to the optimizer perspective, the feedback $\bm{P}_{t}(\cdot)$ in closed-loop control can be either linear and first-order, or nonlinear and higher-order. However, two major challenges arise: (i) adopting nonlinear or high-order optimization techniques, as in TTT, Lattice, or MIRAS, will hinder chunk-wise parallelism; and (ii) recurrent models suffer from the well-known issue of exponential gradient explosion [philipp2017exploding] during training. If we follow DeltaNet using first-order feedback but initialize a new vector [yang2025path] to interact with the state, it becomes difficult to ensure that the spectral radius of the state transition matrix remains below one. Therefore, considering these factors, Comba follows previous models using the $\bm{k}$ vector to interact with the state $\bm{S}$, while introducing a special treatment of feedback strength detailed in the following sections. For a standard Comba with forget gate $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha_{t}$}, state feedback factor $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\beta}_{t}$}, input gate $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}$}, and output feedback factor $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}d$}:
@@
In practice, the scalar gating will be limited to the interval $(0,1)$, and $\|\bm{k}_{t}\|=1$ (L2 Norm). DeltaNet formulates the state transition in an IPLR structure, while RWKV-7 improves expressiveness via a DPLR extension with LoRA-style [hu2022lora] diagonal initialization, i.e., $\operatorname{diag}(\bm{w}_{t})-\beta_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal}$. However, our results indicate that the low-rank form will impair model capacity. Moreover, recent efforts [bick2024transformers, lan2025liger, bick2025llamba, paliotta2025thinking] aim to distill Transformers into recurrent structures to leverage prior knowledge of large-scale pretrained Transformers, where a key design principle is to minimize additional parameters. To this end, Comba adopts an SPLR structure that only introduces a data-dependent scalar, achieving superior empirical performance compared to RWKV-7 (as implemented in FLA [yang2024fla]), along with a 2 $\times$ pretraining acceleration (the memory usage is only half). Recent works [siems2025deltaproduct, grazzi2024unlocking] have also improved IPLR structures by extending their eigenvalues into the negative domain, e.g., $\bm{I}-2\beta_{t}\bm{k}_{t}\bm{k}_{t}^{\intercal}$, to enhance the model’s state-tracking capability. Notably, SPLR naturally admits negative eigenvalues. For fair comparison, in Table 4, we introduce a variant named Comba-$\operatorname{iplr}$. However, such models with global state decay tend to overfit, and the SPLR structure remains optimal and is more structurally aligned with control theory.
@@
Comba introduces additional output feedback, from an optimization perspective, this is equivalent to incorporating a similarity optimization objective $\left\langle\bm{q}_{t},d\bm{k}_{t}\right\rangle$ with factor $d$. In a neural memory perspective, $\bm{k}$ ensures that memory $\bm{v}$ is stored as clearly as possible, enabling precise querying by $\bm{q}$. This optimization objective directly facilitates this process, and significantly reduces the model’s perplexity, thereby enhancing performance (§4.2). Empirically, initializing $d$ to 0.02 improves performance for some smaller models (e.g., 340M), enabling gradual learning of the similarity between $\bm{q}$ and $\bm{k}$. For larger models (e.g., 1.3B, 2.7B), initializing $d$ to 1 leads to the greatest performance [^2] We find that $d=1$ yields better performance in most cases, which is the default configuration for Comba.. While prior research has largely focused on improving gating mechanisms or optimizing state updates via key-value operations, few models have explicitly modified the output pathway (i.e., the query @ State). We think there appears to be a potential connection between Comba and MesaNet [von2025mesanet], as MesaNet also performs query correction in the output stage by solving a closed-form recursive least squares problem [benesty2011recursive], resulting in $\bm{q}_{t}=\left(\bm{H}_{t}+\bm{\Lambda}\right)^{-1}\bm{q}_{t}$. Notably, the correction in Comba can be implemented with a single line of code $\bm{q}_{t}-d\bm{k}_{t}$ and is applicable to nearly all RNN variants. In App. A.2, we also provide an alternative explanation, namely that $d$ is equivalent to the residual matrix $\bm{D}$ in Mamba [gu2023mamba].
@@
Model Forget Gate $\bm{\alpha}_{t}$ Range Input Gate $\bm{\beta}_{t}$ Range GLA [yang2023gated] $\operatorname{sigmoid}(\bm{W}_{1}\bm{W}_{2}\bm{x}_{t})^{\frac{1}{\tau}}\bm{1}^{\intercal}$ (0, 1) N/A 1 Mamba2 [dao2024transformers] $\exp\left(-a\operatorname{softplus}\left(\bm{W}_{\alpha}\bm{x}_{t}+c\right)\right)$ $\sim$ 1 $\operatorname{softplus}\left(\bm{W}_{\alpha}\bm{x}_{t}+c\right)$ $\sim 0$ MetaLA [chou2024metala] $\operatorname{sigmoid}(\bm{W}_{\alpha}\bm{x}_{t})\bm{1}^{\intercal}$ (0, 1) $\bm{1}-\operatorname{sigmoid}(\bm{W}_{\alpha}\bm{x}_{t})$ (0, 1) Comba (ours) $\exp\left(-a\operatorname{softplus}\left(\bm{W}_{\alpha}\bm{x}_{t}+c\right)\right)$ $\sim$ 1 $\operatorname{sigmoid}(\bm{W}_{\beta}\bm{x}_{t})$ $\tilde{\beta}_{t}$ << $\beta_{t}$ $\in$ (0, 1)